# Detecting and Interpreting NSFW Prompts in Text-to-Image Models through Uncovering Harmful Semantics

### v2

@@ A.9 Multi-Categories Classifier
Table 7: The evaluation of different categories.
A.10 The Impact of Training Data Size
[…]
Figure 10: The impact of training data size.
Training Data Size TPR FPR ACC F1 Score AUROC AUPRC TPR@FPR 1%
10 0.8971 0.0278 0.9383 0.9074 0.9815 0.9674 0.8295
50 0.9156 0.0164 0.9514 0.9330 0.9880 0.9801 0.8957
100 0.9405 0.0180 0.9566 0.9438 0.9902 0.9846 0.9151
500 0.9624 0.0091 0.9688 0.9678 0.9921 0.9914 0.9629
1000 0.9633 0.0096 0.9687 0.9676 0.9930 0.9915 0.9633
Table 8: The impact of training data size in CLIP-L.
A.11 Implementation Details of Adaptive Attack
A.11.1 Black-Box Adaptive Attack
In the black-box scenario, we assume the attacker has no knowledge of the model’s details but can choose prompts and query the model to obtain output. We integrate HiddenGuard into the text encoder of Stable Diffusion. When a potential NSFW prompt is detected, the model will refuse to generate the image. We select 100 prompts with clear NSFW semantics from NSFW200 as target prompts and use SneakyPrompt to attack the model.
A.11.2 White-Box Adaptive Attack
In the white-box scenario, we assume attackers can only generate images through queries. However, they possess a local copy of the text encoder identical to the target model and are aware of the HiddenGuard defense strategy. The attacker can target HiddenGuard by modifying the loss function to conduct a specific attack. The MMA attack’s loss aims to make the conditional embeddings of the adversarial prompt and the target prompt as similar as possible. To effectively attack HiddenGuard, we incorporate $Score(p)$ as $L_{\text{HiddenGuard}}$ into the original loss function. This transforms the objective of the loss function to minimize the NSFW Score while ensuring that the semantics of the adversarial prompt closely align with the target prompt.:
$L=L_{\text{MMA}}+\lambda\times L_{\text{HiddenGuard}}$ (14)
Where $\lambda$ is a weighting factor that balances between the two components. Based on this foundation, we implement a target truncation strategy. Specifically, once $L_{\text{HiddenGuard}}$ exceeds the threshold by a small margin, we stop optimizing it and shift our primary focus to optimizing $L_{\text{MMA}}$. This enables adversarial prompts to approximate the semantics of the target prompt as closely as possible while avoiding detection. Consequently, the final loss is formulated as:
$L=L_{\text{MMA}}+\lambda\times\max(L_{\text{HiddenGuard}},\tau-\epsilon),$ (15)
Where $\tau$ is the threshold and $\epsilon$ is the margin. When updating the best prompt, we ensure that the $L_{\text{HiddenGuard}}$ surpasses the threshold. We conduct the attack using the default settings of MMA and evaluate it on 100 target prompts.
A.12 More Results of The Ablation Study
Figure 11: Effectiveness of each layer.
Figure 11 presents the results when using each layer for detection. In all three text encoders, even when using attention heads from a single layer, many layers still achieve high accuracy. For CLIP-L and CLIP-G, the middle layers tend to have higher accuracy. Conversely, in the T5 model, the later layers exhibit higher accuracy. This highlights the distinct characteristics of the two types of text encoders. Although using attention heads from a single layer can achieve high accuracy, we recommend using the original HiddenGuard method for the highest precision in detection.
In Tables 9 through 11, we present more detailed results from the ablation study. The conclusions derived from these tables are broadly consistent with those from Figure 11. In all three text encoders, even when using attention heads from a single layer, many layers still achieve high accuracy. For CLIP-L and CLIP-G, the middle layers tend to have higher accuracy, while the early and final layers show lower accuracy. Conversely, in the T5 model, the later layers exhibit higher accuracy. This highlights the distinct characteristics of the two types of text encoders.
Figure 12: PCA maps of all attention heads.
A.13 Discussion
@@
Practicality. Previous research (Qi et al., 2024) has identified key qualities for a safe and secure generative model, including integrity, robustness, alignment, and interpretability. Building on this foundation, we propose that a practical defense framework should adhere to the following principles: First, the defense must be integrated, offering protection against both conventional threats and adversarial attacks. Second, the defense should remain robust amidst changing external conditions, addressing issues such as distribution shifts and adaptive attacks. Third, the model with the defensive mechanism should align with the original model, preserving its effectiveness and efficiency. Lastly, the defense should be interpretable, enabling the timely identification of anomalous behaviors to prevent potential hazards.
@@
HiddenGuard excels in these four areas compared to prior methods. Firstly, HiddenGuard maintains an accuracy rate exceeding 95% across various models, significantly outperforming earlier approaches. Secondly, it requires only a small number of samples for data augmentation, simplifying updates. Thirdly, its classification process is transparent, and it provides interpretative tools to help users understand target prompts. Lastly, HiddenGuard does not impact image generation quality and incurs only negligible computational overhead. These all demonstrate its practicality.
@@
Limitations and Future Work. HiddenGuard extracts NSFW features from hidden states, showcasing the potential for extracting specific concepts from the text encoder. However, this study does not explore the extraction of other concepts, which warrants further investigation.
@@
Due to its low computational cost, HiddenGuard utilizes all attention heads for detection. However, to better understand the unique characteristics of text encoders like CLIP, it is essential to analyze the semantic focus of individual attention heads. Exploring the roles and properties of each attention head remains a valuable direction for future research.
@@
Despite HiddenGuard demonstrating robust capabilities, no defensive method is foolproof. Therefore, in practical applications, it should be combined with other techniques, such as image moderation. Given that CLIP is a multimodal model, there is potential for adapting HiddenGuard to the image domain to achieve more powerful post-generation defenses.
Layer TPR FPR ACC F1 Score AUROC AUPRC TPR@FPR 1%
1 1.0000 0.9723 0.2730 0.4097 0.9831 0.9581 0.7328
2 0.9982 0.4935 0.6305 0.5768 0.9852 0.9645 0.7835
3 0.9989 0.5949 0.5549 0.5310 0.9870 0.9692 0.8148
4 0.9813 0.1049 0.9168 0.8561 0.9906 0.9773 0.8383
5 0.9571 0.0411 0.9585 0.9208 0.9923 0.9810 0.8598
6 0.9861 0.0824 0.9348 0.8842 0.9940 0.9852 0.8984
7 0.9559 0.0251 0.9701 0.9416 0.9947 0.9871 0.9135
8 0.9653 0.0332 0.9664 0.9355 0.9945 0.9872 0.9103
9 0.9906 0.0983 0.9241 0.8682 0.9953 0.9887 0.9230
10 0.9619 0.0166 0.9780 0.9566 0.9968 0.9918 0.9393
11 0.9187 0.0079 0.9736 0.9460 0.9961 0.9901 0.9317
12 0.8605 0.0031 0.9625 0.9205 0.9968 0.9917 0.9374
13 0.8888 0.0034 0.9694 0.9361 0.9975 0.9934 0.9459
14 0.9381 0.0090 0.9776 0.9549 0.9971 0.9925 0.9402
15 0.9731 0.0210 0.9775 0.9562 0.9972 0.9929 0.9482
16 0.9358 0.0067 0.9788 0.9570 0.9977 0.9937 0.9573
17 0.9646 0.0163 0.9789 0.9584 0.9969 0.9918 0.9402
18 0.9708 0.0176 0.9795 0.9598 0.9972 0.9924 0.9431
19 0.9635 0.0143 0.9801 0.9606 0.9973 0.9929 0.9511
20 0.9639 0.0180 0.9775 0.9557 0.9967 0.9916 0.9381
21 0.9555 0.0179 0.9754 0.9514 0.9965 0.9905 0.9324
22 0.9694 0.0250 0.9736 0.9488 0.9963 0.9901 0.9240
23 0.9769 0.0359 0.9673 0.9378 0.9959 0.9888 0.9123
24 0.9726 0.0207 0.9776 0.9563 0.9971 0.9925 0.9438
Table 9: Effectiveness of each layer in T5.
Layer TPR FPR ACC F1 Score AUROC AUPRC TPR@FPR 1%
1 0.8580 0.0055 0.9601 0.9156 0.9914 0.9814 0.8860
2 0.6221 0.0000 0.9046 0.7670 0.9965 0.9923 0.9523
3 0.8646 0.0006 0.9654 0.9265 0.9979 0.9953 0.9733
4 0.8223 0.0007 0.9547 0.9015 0.9977 0.9949 0.9719
5 0.9201 0.0025 0.9779 0.9546 0.9978 0.9950 0.9710
6 0.9299 0.0030 0.9801 0.9592 0.9980 0.9955 0.9710
7 0.9452 0.0049 0.9825 0.9647 0.9979 0.9952 0.9678
8 0.9472 0.0063 0.9820 0.9636 0.9975 0.9939 0.9651
9 0.9169 0.0023 0.9773 0.9532 0.9981 0.9954 0.9715
10 0.9548 0.0069 0.9834 0.9667 0.9979 0.9949 0.9678
11 0.9644 0.0225 0.9742 0.9496 0.9958 0.9893 0.9258
12 0.8497 0.0040 0.9591 0.9129 0.9940 0.9856 0.9039
Table 10: Effectiveness of each layer in CLIP-L.
Layer TPR FPR ACC F1 Score AUROC AUPRC TPR@FPR 1%
1 0.9873 0.4277 0.6889 0.6409 0.9747 0.9571 0.7752
2 0.8819 0.0388 0.9389 0.8903 0.9780 0.9601 0.7696
3 0.9484 0.0389 0.9575 0.9261 0.9895 0.9812 0.8969
4 0.7610 0.0022 0.9313 0.8616 0.9869 0.9767 0.8629
5 0.8246 0.0022 0.9491 0.9011 0.9943 0.9891 0.9285
6 0.7750 0.0014 0.9358 0.8715 0.9922 0.9851 0.8977
7 0.7512 0.0006 0.9297 0.8572 0.9944 0.9895 0.9252
8 0.6460 0.0002 0.9003 0.7847 0.9937 0.9878 0.9166
9 0.6608 0.0003 0.9044 0.7954 0.9931 0.9865 0.9067
10 0.6563 0.0002 0.9032 0.7922 0.9931 0.9864 0.9047
11 0.8272 0.0032 0.9491 0.9014 0.9929 0.9860 0.8961
12 0.8188 0.0011 0.9483 0.8990 0.9946 0.9902 0.9299
13 0.8666 0.0037 0.9598 0.9239 0.9954 0.9907 0.9270
14 0.7846 0.0014 0.9385 0.8776 0.9935 0.9878 0.9227
15 0.9061 0.0051 0.9699 0.9443 0.9948 0.9906 0.9399
16 0.9625 0.0187 0.9760 0.9576 0.9948 0.9908 0.9365
17 0.8793 0.0046 0.9628 0.9300 0.9940 0.9890 0.9235
18 0.8162 0.0178 0.9355 0.8768 0.9808 0.9613 0.7379
19 0.9334 0.0095 0.9744 0.9535 0.9948 0.9903 0.9346
20 0.8387 0.0020 0.9532 0.9097 0.9946 0.9900 0.9320
21 0.8145 0.0014 0.9468 0.8960 0.9940 0.9889 0.9293
22 0.8131 0.0023 0.9458 0.8940 0.9944 0.9891 0.9162
23 0.8949 0.0055 0.9665 0.9375 0.9944 0.9893 0.9346
24 0.9457 0.0155 0.9736 0.9527 0.9940 0.9884 0.9252
25 0.8563 0.0037 0.9570 0.9179 0.9940 0.9882 0.9213
26 0.5157 0.0031 0.8616 0.6769 0.9873 0.9690 0.7483
27 0.9297 0.0140 0.9702 0.9460 0.9940 0.9874 0.9092
28 0.7557 0.0014 0.9303 0.8590 0.9944 0.9886 0.9190
29 0.8971 0.0115 0.9628 0.9313 0.9929 0.9852 0.8766
30 0.5235 0.0296 0.8448 0.6547 0.9124 0.8260 0.3167
31 0.8635 0.0034 0.9592 0.9224 0.9939 0.9889 0.9289
32 0.7670 0.0067 0.9297 0.8598 0.9877 0.9753 0.8159
Table 11: Effectiveness of each layer in CLIP-G.

### v3

@@ V-C Red-Teaming Data Augmentation
In the practical application of AEIOU, we can also collect adversarial prompts that successfully bypass defenses through manual screening. These prompts can then be added to the training dataset with appropriate weighting, allowing for continuous updates of the defense model.
VI Interpretation Experiments
In this section, we validate AEIOU’s interpretability. This not only makes AEIOU’s classification more transparent and trustworthy but also aids users in further understanding the semantics of prompts. We assess the reliability of interpretations across both text and image modalities.
VI-A Text-Based Interpretation
In the text modality, our interpretation aids users in understanding the semantics of a prompt by identifying tokens containing NSFW semantics. After obtaining the interpretation result $E(p)_{i}$ for each token in the prompt $p$, we sequentially remove the corresponding tokens from $p$ in descending order of $E(p)_{i}$ and observe the changes in the NSFW score. We compare our interpretation method against two baselines. The first, a random-based method, removes tokens from the prompt at random. The second, an embedding-based method, uses the CLIP score to evaluate the semantic similarity of each token to the term ”NSFW” and removes tokens in descending order of this similarity. The experimental results are shown in Figure 8. Compared to both baselines, removing tokens based on our interpretation method more rapidly reduces the NSFW semantic content of the prompt. This demonstrates the effectiveness of AEIOU in both explaining NSFW prompts and locating the specific tokens responsible for the NSFW content.
Figure 8: Text-based interpretation.
Additionally, we conduct further experiments on image generation using the 100 prompts mentioned above. We compare the images generated from the original prompt with those generated from the prompt after removing the NSFW token. For original prompts, 87 out of the 100 generated images are NSFW. After removing NSFW tokens, only 5 prompts result in NSFW images. This further demonstrates the effectiveness of our interpretation method.
VI-B Image-Based Interpretation
Directly removing tokens with NSFW semantics can prevent the generation of NSFW images. However, it significantly disrupts the original intent of the prompt. To address this drawback, we manipulate the embeddings of each token in the image-based interpretation process, gradually eliminating NSFW semantics and observing changes in the generated images. We use the same 100 prompts as in the previous section. We evaluate the degree to which the generated images contain NSFW semantics and how closely they align with the original prompt semantics.
Figure 9 illustrates the variation in the number of NSFW images generated and the semantic similarity between the images and prompts as the parameter $\beta$ changes. The semantic similarity is evaluated using the CLIP Score. As $\beta$ increases, the number of NSFW images decreases progressively. Although the CLIP Score also shows a decreasing trend, the overall deviation from the original image’s CLIP Score remains minimal. This indicates that the NSFW semantics of the prompt are effectively mitigated while preserving other semantic information as much as possible.
However, because this method directly modifies semantics within the hidden states, the resulting conditional embeddings deviate from the normal distribution, often leading to lower-quality images. Therefore, this approach is supposed to aid in understanding the representation of NSFW semantics within the text encoder. It cannot be directly applied to erase specific concepts from images.
VII Discussion
@@
Practicality. Previous research [qi2024ai] has identified key qualities for a safe and secure generative model, including integrity, robustness, alignment, and interpretability. Building on this foundation, we propose that a practical defense framework should adhere to the following principles: First, the defense must be integrated, offering protection against both conventional threats and adversarial attacks. Second, the defense should remain robust amidst changing external conditions, addressing issues such as distribution shifts and adaptive attacks. Third, the model with the defensive mechanism should align with the original model, preserving its effectiveness and efficiency. Lastly, the defense should be interpretable, enabling the timely identification of anomalous behaviors to prevent potential hazards.
@@
AEIOU excels in these four areas compared to prior methods. Firstly, AEIOU maintains an accuracy rate exceeding 95% across various models, significantly outperforming earlier approaches. Secondly, it requires only a small number of samples for data augmentation, simplifying updates. Thirdly, its classification process is transparent, and it provides interpretative tools to help users understand target prompts. Lastly, AEIOU does not impact image generation quality and incurs only negligible computational overhead. These all demonstrate its practicality.
Figure 9: Image-based interpretation.
@@
Limitations and Future Work. AEIOU extracts NSFW features from hidden states, showcasing the potential for extracting specific concepts from the text encoder. However, this study does not explore the extraction of other concepts, which warrants further investigation.
@@
Due to its low computational cost, AEIOU utilizes all attention heads for detection. However, to better understand the unique characteristics of text encoders like CLIP, it is essential to analyze the semantic focus of individual attention heads. Exploring the roles and properties of each attention head remains a valuable direction for future research.
@@
Despite AEIOU demonstrating robust capabilities, no defensive method is foolproof. Therefore, in practical applications, it should be combined with other techniques, such as image moderation. Given that CLIP is a multimodal model, there is potential for adapting AEIOU to the image domain to achieve more powerful post-generation defenses.
VIII Conclusion
In this paper, we propose a unified defense framework against NSFW prompts in T2I models named AEIOU. The AEIOU framework, being adaptable, efficient, interpretable, optimizable, and unified, has demonstrated superior performance that far exceeds previous defense methods. In addition to detection, we provide interpretability methods to help understand the semantics of NSFW prompts and the generation process of NSFW images. Experimental results show that AEIOU exhibits strong capabilities in defending against both normal attacks and adaptive attacks.